OpenStudy (tanabugg):

3x + 4(x-6) - 3(x - 7) 4x- 45 4x -3 10x - 3 10x +45

pooja195 (pooja195):

Can you attempt to do the problem ? :)

pooja195 (pooja195):

Show your steps

OpenStudy (tanabugg):

yes

OpenStudy (tanabugg):

So, first would I combine (x - 6) and (x - 7) right?

pooja195 (pooja195):

Nope don't over think this :-) Start by distributing

OpenStudy (prettygirl_shynice):

srry

OpenStudy (landon34):

YAY guys im cleaned of the spyware!

OpenStudy (prettygirl_shynice):

i dint mean to be mean

pooja195 (pooja195):



OpenStudy (landon34):

do you know how to do it @Prettygirl_shynice ?

OpenStudy (landon34):

3x + 4(x + -6) + -3(x + -7) = 0 Reorder the terms: 3x + 4(-6 + x) + -3(x + -7) = 0 3x + (-6 * 4 + x * 4) + -3(x + -7) = 0 3x + (-24 + 4x) + -3(x + -7) = 0 Reorder the terms: 3x + -24 + 4x + -3(-7 + x) = 0 3x + -24 + 4x + (-7 * -3 + x * -3) = 0 3x + -24 + 4x + (21 + -3x) = 0 Reorder the terms: -24 + 21 + 3x + 4x + -3x = 0 Combine like terms: -24 + 21 = -3 -3 + 3x + 4x + -3x = 0 Combine like terms: 3x + 4x = 7x -3 + 7x + -3x = 0 Combine like terms: 7x + -3x = 4x -3 + 4x = 0 Solving -3 + 4x = 0 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '3' to each side of the equation. -3 + 3 + 4x = 0 + 3 Combine like terms: -3 + 3 = 0 0 + 4x = 0 + 3 4x = 0 + 3 Combine like terms: 0 + 3 = 3 4x = 3 Divide each side by '4'. x = 0.75 Simplifying x = 0.75

OpenStudy (prettygirl_shynice):

omm i do im scraed

OpenStudy (landon34):

dont be scared. if you dont know it, tell him.